# Non-abelian Rees construction and pure motives

### v1

@@ 2.3.2. Remarks
1) This is a Galois-type correspondence. In classical Galois theory, only groups and homogeneous spaces occur on the group-theoretic side of the dictionary. The first appearance of quasi-homogeneous spaces arose in differential Galois theory (beyond the differential field case), where an anti-equivalence between quasi-homogeneous varieties and certain ‘solution algebras’ was established in [5], and advantage was taken of the rich theory of quasi-homogeneous varieties, cf. [7]. The notion of ‘solution algebras’ and this anti-equivalence were later recast in a general monoidal setting [33].
2) $H$ may not be reductive but is an observable subgroup of $G$, since ${\rm{X}}_{0}=G/H$ is quasi-affine. An equivalent characterization of observability is: any $V\in{\rm{Rep}}\,H$ is the restriction of some $W\in{\rm{Rep}}\,G$, cf. [21].
@@ 2.3.4. Example:
More generally, if ${\mathcal{C}}={\rm{Rep}}\,H$ for a linear algebraic group $H$, then $\mathcal{C}$} satisfies $F1)$ if and only if the unipotent radical of $H$ is either trivial (in which case $\rm X$ is a point) or $\mathbb{G}_{a}$ (in which case $\rm X$ is $\mathbb{A}^{2}$) [35].
@@ 2.3.5. Example: multifiltrations
only provides a lax tensor functor, which does not preserve duals141414a lax tensor functor which preserve duals is a tensor functor [4, App.]..
@@
$(\ast)$ there is a basis $v_{i}$ of homogeneous elements of ${\rm{Gr}}\,\underline{V}=\oplus_{{\bf n}\in\mathbb{N}^{n}}\,{\rm{Gr}}_{\bf n}\underline{V}$, and liftings $\tilde{v}_{i}$ of $v_{i}$ which form a basis of $V$.
@@
Moreover, under $(\ast)$, the $n$-variable Rees construction provides ${\mathbb{G}}_{m}^{n}$-equivariant vector bundles over ${\mathbb{A}}^{n}$ [49, 2.2], and ${\mathcal{C}}\to{\rm{Rep}}\,\mathbb{G}_{m}^{n}$ factors through ${\rm{Vec}}_{\mathbb{G}_{m}^{n}}\mathbb{A}^{n}$; it is full, hence factors through $\bar{\mathcal{C}}$ [6, 1.4.7]. Therefore there is a quotient torus $G$ of $\mathbb{G}_{m}^{n}$ such that $\bar{\mathcal{C}}\cong{\rm{Rep}}\,G$. If $G$ is a split torus, there is an equivariant map ${\mathbb{A}}^{n}\to{\rm X}$, which sends the origin to the fixed point $o$. The classical $n$-dimensional Rees construction is thus closely related, but not necessarily equal to O’Sullivan’s version.
Conversely, given a split torus $G$ over $K$ and a fix-pointed affine toric variety $\rm X$, the tensor category ${\rm{Vec}}_{G}{\rm X}$ is equivalent to the pseudo-abelian rigid full tensor subcategory of $n{\rm{FilVec}}_{K}$ generated by $V=X^{\ast}(G)\otimes K$ endowed with $n$ filtrations attached to each face (not compatible in the sense of $(\ast)$, unless the cone is simplicial, but compatible in the sense of Klyachko cf. [26][39, 2.3, 2.4]).
@@ 2.3.6. Example: the “Fibonacci category"
Let $\mathcal{C}$} be the rigid tensor category generated by two objects of rank one $L_{1},L_{2}$, with one non-zero morphism $L_{1}^{a}\to L_{2}^{b}$ (up to scaling) for $a,b\in\mathbb{N}$, if and only if $b/a\leq\frac{1+\sqrt{5}}{2}$.
@@ 3. Properties of the restricted non-abelian Rees construction
Let $\mathcal{C}$} be again an even $K$-tensor category (1.2.3), with a fiber functor $\omega:{\mathcal{C}}\to{\rm{Vec}}_{K}$, a tensor section $\sigma$ of $\pi:{\mathcal{C}}\to\bar{\mathcal{C}}$, $H\subset G$ the tannakian groups attached to $({\mathcal{C}}_{tann},\omega)$ and $(\bar{\mathcal{C}},\bar{\omega}=\omega\sigma)$ respectively.
@@ 3.1. A formula for ${\mathcal{O}}({\rm{X}})$
${\mathcal{O}}({\rm{X}})=\oplus_{{\rm{indec}}\,\mathcal{V}}\;{\mathcal{C}}(\mathcal{V},{\bf 1})\otimes V\,\in{\rm{REP}}\,G,$
where $\mathcal{V}$ runs over the isomorphism classes of indecomposable objects, $V=\omega({\mathcal{V}})$ denotes the corresponding irreducible $G$-representation, and ${\mathcal{O}}({\rm{X}})$ is a subalgebra of ${\mathcal{O}}({\rm{G}})=\oplus_{{\rm{indec}}\,\mathcal{V}}\;\omega(\mathcal{V}^{\vee})\otimes V\,\in{\rm{REP}}\,G$ via the $G$-action on $V$. This is an immediate consequence of O’Sullivan’s construction and Schur’s lemma.
@@
Indeed, ${\mathcal{C}}({\mathcal{W}},{\bf 1})=Hom_{G}(W,{\mathcal{O}}({\rm{G}}))$ and the formula together with Schur’s lemma $\dim Hom_{G}(W,V)\leq 1$ gives the result.
@@ 3.2. Functorial properties
Here $\mathcal{C}$} and $\mathcal{C}$$^{\prime}$ denote even $K$-tensor categories as above (satisfying P1, P2), and we consider tensor functors ${\mathcal{C}}\to{\mathcal{C}}^{\prime}$.
@@ 3.2.1. Remarks
2) Let us consider again the abelian tensor hull $T({\mathcal{C}})$ (2.1.4). The (absolutely flat) $K$-algebra $T({\mathcal{C}})({\bf 1},{\bf 1})$ may be huge. In [24, 6.6], Kahn constructed a spectral map $\operatorname{Spec}\,T({\mathcal{C}})({\bf 1},{\bf 1})\to\operatorname{Spec}^{\otimes}{\mathcal{C}}$, which in our case (by the previous remark) is a homeomorphism if one gives $\operatorname{Spec}^{\otimes}{\mathcal{C}}$ the constructible topology. Moreover, by Theorem 2.3.1 c), the points of $\operatorname{Spec}^{\otimes}{\mathcal{C}}$ are in bijection with the $G$-orbits in $\rm X$. Whether there are finitely many $G$-orbits (thus - equivalently - whether $T({\mathcal{C}})({\bf 1},{\bf 1})$ is finite) is well-understood and closely related to sphericity, cf. [7, §4]. There may be infinitely many orbits, but there are only finitely many if $\rm X_{0}$ is spherical.
The map $\rm X\to\operatorname{Spec}^{\otimes}{\mathcal{C}}$ which sends $x\in\rm X$ to the kernel of the tensor functor $\omega_{x}$ is not continuous, but become continuous (an epimorphism of profinite spaces, if fact the quotient by the $G(K)$-action), if one endows source and target with the constructible topology.
@@
In the case of the Fibonacci category $\mathcal{C}$} (2.3.6), viewed of the filtered union of subcategories ${\mathcal{C}}_{(n)}$ satisfying P3, each $\operatorname{Spec}^{\otimes}{\mathcal{C}}_{(n)}$ consists of 4 points (corresponding to the facets of the cone of the toric surface ${\rm X}_{(n)}$), while $\operatorname{Spec}^{\otimes}{\mathcal{C}}$ consists only of 3 points: the tensor ideals $0,{\mathcal{N}}$ and the one generated by morphisms $L_{1}^{a}\to L_{2}^{b}$ for a,b>0.a,b>0.151515in fact, in this projective system, the 1st, 2nd, and 4th points map to the 1st, 2nd, and 4th points, while the 3rd point maps to the 4th point.
3) $G$-orbits of minimal nonzero dimension in $\rm X$ are of special interest (under condition P3). The normalization $\rm X^{\prime}$ of the closure of such an orbit is still a fix-pointed affine quasi-homogeneous variety: it is a ‘HV variety’, i.e. the closure of the orbit of a highest weight vector in some irreducible $G$-representation. The corresponding group $H^{\prime}$ is a quasi-parabolic subgroup of $G$, i.e. the kernel of a character of infinite order if some parabolic subgroup of $G$, cf. [7, §1, 3.5].
Moreover, if $\rm X$ is irreducible, there is a surjective equivariant map from $\rm X$ to such a $\rm X^{\prime}$ (Bogomolov), cf. e.g. [21, 7.8].
@@ 3.3. Extension of scalars
Let $K^{\prime}/K$ be a finite extension. Let ${\mathcal{C}}_{K^{\prime}}^{\natural}$ be the pseudo-abelian closure of the naive extension of scalars of $\mathcal{C}$} (which satisfies the conditions of Theorem 2.3.1 if $\mathcal{C}$} itself satisfies them). Then if the above correspondence over $K$ associates $(G,{\rm X})$ to $\mathcal{C}$}, it associates $(G_{K^{\prime}},{\rm X}_{K^{\prime}})$ to ${\mathcal{C}}_{K^{\prime}}^{\natural}$ over $K^{\prime}$.
@@ 3.4. Criteria of semisimplicity for $\mathcal{C}$}
In some special situations, arguments from the theory of quasi-homogeneous varieties can lead to the conclusion that $\rm X$ must be a point. We give here two such instances, the first one relying on a theorem of Luna, the second one on descent. These criteria do not mention the deformation space $\rm X$ in their statement, but use it in their proof.
@@
Let $K=\bar{K}$. Let $N_{G}(H)$ (resp. $Z_{G}(H)$) be the normalizer (resp. centralizer) of $H$ in $G$, and $W(H):=N_{G}(H)/H$. If $H$ is reductive, which amounts to say that $\rm X_{0}$ is affine (Matsushima), then $N_{G}(H)$ are $Z_{G}(H)$ are reductive, and $W(H)$ is commensurable to $Z_{G}(H)/(Z_{G}(H)\cap H)$, cf. [7, 3.1]. One has $Aut_{G}{\rm X}\subset Aut_{G}{\rm X}_{0}=W(H)$, cf. [7, 4.3] .
If there is a reductive subgroup $H^{\prime}\subset H$ such that $W(H^{\prime})$ is finite, then $H$ is reductive and $\rm X_{0}$ is the unique affine quasi-homogeneous variety containing $\rm X_{0}$ (Luna), cf. [7, 3.1]. Since $\rm X$ is fix-pointed, this forces $\rm X$ to be a point.
@@
Assume $K=\bar{K}$ and let ${\mathcal{C}}=\langle{\mathcal{V}}\rangle$ be such that:
@@
One may replace by an indecomposable factor. Then $\omega(\mathcal{V})$ is an irreducible $G$-representation. By Schur, $Z_{G}(H^{\prime})=K^{*}$.
@@
Let $\omega^{\prime}:{\mathcal{C}}^{\prime}\to{\rm{Vec}}_{K^{\prime}}$ be the extension of $\omega,\,G^{\prime}=G_{K^{\prime}}$ (cf. 3.3).
One has the formula $\mathcal{O}({\rm{X}}^{\prime})=\oplus_{\chi\in X^{*}(G^{\prime})}\,{\mathcal{C}}^{\prime}({\bf 1},V(\chi^{-1}))\otimes\omega^{\prime}(V(\chi))\,$
@@
Now, $\iota$ acts functorially on $\mathcal{C}$$^{\prime}$, $\iota(V(\chi))=V(\iota(\chi))$, thus $\Xi=\iota(\Xi)$. Hence $\Xi=-\Xi$ is a subgroup of $X^{*}(G^{\prime})$, so that ${\rm{X}}^{\prime}=G^{\prime}/H^{\prime}$. But $\rm X$$^{\prime}$ is fix-pointed, hence is a point, and so is $\rm X$. $\square$
@@ 4.1. General setting
Grothendieck motives (for homological equivalence) form a pseudo-abelian rigid tensor category. It is built from projective varieties over a field $k$, morphisms being algebraic correspondences (of degree 0) modulo homological equivalence. In the sequel, we consider only correspondences with coefficients in $K$, the field of coefficients of the given cohomology theory.
@@
Any tensor subcategory $\mathcal{C}$} of motives cut on products of finitely many varieties $Y_{m}$ satisfies by definition condition P2 of 2.2. In order to have P1 (embedding into ${\rm{Vec}}_{K}$), one needs to twist the commutativity constraint according to the Koszul rule, assuming the sign conjecture for the $Y_{m}$, i.e. that the Künneth projectors on the even part of cohomology are algebraic, so that $\mathcal{C}$} becomes even161616under Voevodsky’s smash-nilpotence conjecture, one could perform this at the level of non-integral tensor category of Chow motives; then O’Sullivan’s construction yields a $G$-scheme $\rm X$ which is a ‘fat point’ (the reduced scheme is a point), cf. [2, §2]..
@@
Here $\bar{\mathcal{C}}$ is the category for numerical equivalence cut on products of $Y_{m}$’s, and $G$ (the numerical motivic Galois group) is the tannakian group of $\bar{\mathcal{C}}$ with respect to the fiber functor given by $\bar{\omega}=\omega\sigma$, where $\sigma$ is a fixed tensor section of ${\mathcal{C}}\to\bar{\mathcal{C}}$ and $\omega$ is the fiber functor on $\mathcal{C}$} given by the selected cohomology.
According to Theorem 2.3.1, the deformation space $\rm X$} which interpolates between $\bar{\mathcal{C}}$ and $\mathcal{C}$} is the limit of a tower of fix-pointed affine quasi-homogeneous varieties with the same $G$ and same open orbit $G/H$.
@@
What is $H$?
If the Künneth projectors are algebraic and char $k=0$, $H$ is the “motivated" motivic Galois group [1], the tannakian group of the tensor category ${\mathcal{C}}_{mot}$ obtained from $\mathcal{C}$} by adding formally the inverses of Lefschetz operators $L_{Y,i}$ (the isomorphisms $H^{i}(Y)\to H^{2d-i}(Y)(d-i)$ given by the (di)(d-i)th iterated cup-product with the class of an ample line bundle on $Y$).
Indeed, ${\mathcal{C}}_{mot}$ is abelian, hence a target from ${\mathcal{C}}_{tann}$; and $L_{Y,i}$, which induces a morphism in ${\mathcal{C}}_{tann}$ with trivial kernel and cokernel, is an isomorphism - from which it is easy to conclude that ${\mathcal{C}}_{mot}={\mathcal{C}}_{tann}$. Moreover, ${\mathcal{C}}_{mot}$ is semisimple [1], hence $H$ is a reductive subgroup of $G$.
In this context, the caveat 2.1.3 takes the form that it is not clear whether the notion of a motivated class on $Y$ is independent of the set of auxiliary varieties $Y_{i}$ on which it is modelled.
@@
for $\mathcal{C}$} is equivalent to
@@
The deformation space $\rm X$} thus appears as an obstruction to conjecture D. Closed $G$-subvarieties in $\rm X$ correspond to adequate equivalence relations intermediate between homological equivalence and numerical equivalence.

### v2

@@ 2.3.2. Remarks
1) This is a Galois-type correspondence. In classical Galois theory, only groups and homogeneous spaces occur on the group-theoretic side of the dictionary. The first appearance of quasi-homogeneous spaces arose in differential Galois theory (beyond the differential field case), where an anti-equivalence between affine quasi-homogeneous varieties and certain ‘solution algebras’ was established in [5], and advantage was taken of the rich theory of affine quasi-homogeneous varieties, cf. [7]. The notion of ‘solution algebras’ and this anti-equivalence were later recast in a general monoidal setting [35].
2) $H$ may not be reductive but is an observable subgroup of $G$, since ${\rm{X}}_{0}=G/H$ is quasi-affine. An equivalent characterization of observability’ is: any $V\in{\rm{Rep}}\,H$ is the restriction of some $W\in{\rm{Rep}}\,G$, cf. [22].
@@ 2.3.4. Example:
More generally, if ${\mathcal{C}}={\rm{Rep}}\,H$ for some linear algebraic group $H$, then $\mathcal{C}$} satisfies P1 if and only if the unipotent radical of $H$ is either trivial (in which case $\rm X$ is a point) or $\mathbb{G}_{a}$ (in which case $\rm X$ is $\mathbb{A}^{2}$) [37].
@@ 2.3.5. Example: multifiltrations
only provides a lax tensor functor which does not preserve duals (a lax tensor functor which preserves duals is a tensor functor [4, App.]).
@@
$(\ast)$ there is a basis $v_{i}$ of homogeneous elements of ${\rm{Gr}}\,\underline{V}=\oplus_{{\bf n}\in\mathbb{N}^{n}}\,{\rm{Gr}}_{\bf n}\underline{V}$, and liftings $\tilde{v}_{i}$ of $v_{i}$ that form a basis of $V$.
@@
Moreover, under $(\ast)$, the $n$-variable Rees construction provides ${\mathbb{G}}_{m}^{n}$-equivariant vector bundles over ${\mathbb{A}}^{n}$ [51, 2.2], and ${\mathcal{C}}\to{\rm{Rep}}\,\mathbb{G}_{m}^{n}$ factors through ${\rm{Vec}}_{\mathbb{G}_{m}^{n}}\mathbb{A}^{n}$; it is full, hence factors through $\bar{\mathcal{C}}$ [6, 1.4.7]. Therefore there is a quotient torus $G$ of $\mathbb{G}_{m}^{n}$ such that $\bar{\mathcal{C}}\cong{\rm{Rep}}\,G$. If $G$ is a split torus, there is an equivariant map ${\mathbb{A}}^{n}\to{\rm X}$, which sends the origin to the fixed point $o$. The classical $n$-dimensional Rees construction is thus closely related, but not necessarily equal to O’Sullivan’s version.
Conversely, given a split torus $G$ over $K$ and a fix-pointed affine toric variety $\rm X$, the tensor category ${\rm{Vec}}_{G}({\rm X})$ is equivalent to the pseudo-abelian rigid full tensor subcategory of $n{\rm{FilVec}}_{K}$ generated by $V=X^{\ast}(G)\otimes K$ endowed with $n$ filtrations attached to each face (not compatible in the sense of $(\ast)$, unless the cone is simplicial, but compatible in the sense of Klyachko cf. [27][41, 2.3, 2.4]).
@@ 2.3.6. Example: the “Fibonacci category"
Let $\mathcal{C}$} be the rigid tensor category generated by two objects of rank one $L_{1},L_{2}$, with one (and only one up to scaling) non-zero morphism $L_{1}^{a}\to L_{2}^{b}\;((a,b)\in\mathbb{N}^{2})$ if and only if $b/a\leq\frac{1+\sqrt{5}}{2}$.
@@ 3. Properties of the restricted non-abelian Rees construction
Let $\mathcal{C}$} be again an even $K$-tensor category (1.2.3), endowed with a fiber functor $\omega:{\mathcal{C}}\to{\rm{Vec}}_{K}$, and a tensor section $\sigma$ of $\pi:{\mathcal{C}}\to\bar{\mathcal{C}}$. Let $H\subset G$ be the tannakian groups attached to $({\mathcal{C}}_{tann},\omega)$ and $(\bar{\mathcal{C}},\bar{\omega}=\omega\sigma)$ respectively.
@@ 3.1. A formula for ${\mathcal{O}}({\rm{X}})$
${\mathcal{O}}({\rm{X}})=\oplus_{{\rm{indec}}\,\mathcal{V}}\;{\mathcal{C}}(\mathcal{V},{\bf 1})\otimes V\,\in{\rm{REP}}\,G.$
Here $\mathcal{V}$ runs over the isomorphism classes of indecomposable objects, $V=\omega({\mathcal{V}})$ denotes the corresponding irreducible $G$-representation, and ${\mathcal{O}}({\rm{X}})$ is a subalgebra of ${\mathcal{O}}({\rm{G}})=\oplus_{{\rm{indec}}\,\mathcal{V}}\;\omega(\mathcal{V}^{\vee})\otimes V\,\in{\rm{REP}}\,G$ via the $G$-action on $V$. This is an immediate consequence of O’Sullivan’s construction and Schur’s lemma.
@@
Indeed, ${\mathcal{C}}({\mathcal{W}},{\bf 1})=Hom_{G}(W,{\mathcal{O}}({\rm{G}}))$ and the formula together with Schur’s lemma ($\dim Hom_{G}(W,V)\leq 1$) gives the result.
@@ 3.2. Functorial properties
Here $\mathcal{C}$} and $\mathcal{C}$$^{\prime}$ denote even $K$-tensor categories as above (satisfying P1, P2). We consider tensor functors ${\mathcal{C}}\stackrel{{\scriptstyle\phi}}{{\to}}{\mathcal{C}}^{\prime}$, and the corresponding equivariant morphisms of quasi-homogeneous varieties ${\rm X}^{\prime}\stackrel{{\scriptstyle f}}{{\to}}{\rm X}$.
@@ 3.2.1. Remarks
2) Let us consider again the abelian tensor hull $T({\mathcal{C}})$ (2.1.4). The (absolutely flat) $K$-algebra $T({\mathcal{C}})({\bf 1},{\bf 1})$ may be huge. In [25, 6.6], Kahn constructed a spectral map $\operatorname{Spec}\,T({\mathcal{C}})({\bf 1},{\bf 1})\to\operatorname{Spec}^{\otimes}{\mathcal{C}}$, which in our case (by the previous remark) is a homeomorphism if one gives $\operatorname{Spec}^{\otimes}{\mathcal{C}}$ the constructible topology. Moreover, by Theorem 2.3.1 c), the points of $\operatorname{Spec}^{\otimes}{\mathcal{C}}$ are in bijection with the $G$-orbits in $\rm X$. Whether there are finitely many $G$-orbits (thus - equivalently - whether $T({\mathcal{C}})({\bf 1},{\bf 1})$ is finite) is well-understood and closely related to sphericity, cf. [7, §4]. There may be infinitely many orbits, but there are only finitely many if $\rm X_{0}$ is spherical.
The map $\rm X\to\operatorname{Spec}^{\otimes}{\mathcal{C}}$ that sends $x\in\rm X$ to the kernel of the tensor functor $\omega_{x}$ is not continuous, but become continuous (an epimorphism of profinite spaces, if fact the quotient by the $G(K)$-action), if one endows source and target with the constructible topology.
@@
In the case of the Fibonacci category $\mathcal{C}$} (2.3.6), viewed of the filtered union of subcategories ${\mathcal{C}}_{(n)}$ satisfying P3, each $\operatorname{Spec}^{\otimes}{\mathcal{C}}_{(n)}$ consists of 4 points (corresponding to the facets of the cone of the toric surface ${\rm X}_{(n)}$), while $\operatorname{Spec}^{\otimes}{\mathcal{C}}$ consists only of 3 points: the tensor ideals $0,{\mathcal{N}}$ and the one generated by morphisms $L_{1}^{a}\to L_{2}^{b}$ for a,b>0.a,b>0.121212in fact, in this projective system of quadruplets ($\operatorname{Spec}^{\otimes}{\mathcal{C}}_{(n)}$), the 1st, 2nd, and 4th points map to the 1st, 2nd, and 4th points, while the 3rd point is sent to the 4th point.
3) $G$-orbits of minimal nonzero dimension in $\rm X$ are of special interest (under condition P3). The normalization $\rm X^{\prime}$ of the closure of such an orbit is still a fix-pointed affine quasi-homogeneous variety: it is the closure of the orbit of a highest weight vector in some irreducible $G$-representation. The corresponding group $H^{\prime}$ is a quasi-parabolic subgroup of $G$, i.e. the kernel of a character of infinite order if some parabolic subgroup of $G$, cf. [7, §1, 3.5].
Moreover, if $\rm X$ is irreducible, there is a surjective equivariant map from $\rm X$ to such a $\rm X^{\prime}$ (Bogomolov), cf. e.g. [22, 7.8].
@@ 3.3. Extension of scalars
Let $K^{\prime}/K$ be a finite extension. Let ${\mathcal{C}}_{K^{\prime}}^{\natural}$ be the pseudo-abelian closure of the naive extension of scalars of $\mathcal{C}$} (which satisfies the conditions of Theorem 2.3.1 if $\mathcal{C}$} itself satisfies them). Then if the correspondence of 2.3.1 over $K$ associates $(G,{\rm X})$ to $\mathcal{C}$}, it associates $(G_{K^{\prime}},{\rm X}_{K^{\prime}})$ to ${\mathcal{C}}_{K^{\prime}}^{\natural}$ over $K^{\prime}$.
@@ 3.4. Criteria of semisimplicity for $\mathcal{C}$}
In some special situations, arguments from the theory of quasi-homogeneous varieties can lead to the conclusion that $\rm X$ must be a point. We give here two such instances, the first one relying on a theorem of Luna [31], the second one on descent. These criteria do not mention the deformation space $\rm X$ in their statement, but use it in their proof.
@@
Let $K=\bar{K}$. Let $N_{G}(H)$ (resp. $Z_{G}(H)$) be the normalizer (resp. centralizer) of $H$ in $G$, and $W(H):=N_{G}(H)/H$. If $H$ is reductive, which amounts to say that the homogeneous space ${\rm X}_{0}=G/H$ is affine (Matsushima’s criterion), then $N_{G}(H)$ are $Z_{G}(H)$ are reductive, and $W(H)$ is commensurable to $Z_{G}(H)/(Z_{G}(H)\cap H)$, cf. [7, 3.1]. One has $Aut_{G}{\rm X}\subset Aut_{G}{\rm X}_{0}=W(H)$, cf. [7, 4.3].
If there is a reductive subgroup $H^{\prime}\subset H$ such that $W(H^{\prime})$ is finite, then $H$ is also reductive and $\rm X_{0}$ is the unique affine quasi-homogeneous variety containing the homogeneous space $\rm X_{0}$ (Luna’s criterion), cf. [7, 3.1]. If $\rm X$ is fix-pointed, this forces $\rm X$ to be a point.
@@
Assume $K=\bar{K}$ and let ${\mathcal{C}}=\langle{\mathcal{V}}\rangle$ (cf. 2.2 for the notation) be such that:
@@
One may replace $\mathcal{V}$ by an indecomposable factor. Then $\omega(\mathcal{V})$ is an irreducible $G$-representation. By Schur’s lemma, $Z_{G}(H^{\prime})=K^{*}$.
@@
Let $\omega^{\prime}:{\mathcal{C}}^{\prime}\to{\rm{Vec}}_{K^{\prime}}$ be the extension of $\omega$, so that $G^{\prime}=G_{K^{\prime}}$ (cf. 3.3). One has
$\mathcal{O}({\rm{X}}^{\prime})=\oplus_{\chi\in X^{*}(G^{\prime})}\,{\mathcal{C}}^{\prime}({\bf 1},V(\chi^{-1}))\otimes\omega^{\prime}(V(\chi))\,$
@@
Now, $\iota$ acts functorially on $\mathcal{C}$$^{\prime}$, $\iota(V(\chi))=V(\iota(\chi))$, thus $\Xi=\iota(\Xi)$. Hence $\Xi=-\Xi$ is a subgroup of $X^{*}(G^{\prime})$, so that ${\rm{X}}^{\prime}=G^{\prime}/H^{\prime}$. But $\rm X$$^{\prime}$ is fix-pointed, hence is a point, therefore so is $\rm X$. $\square$
@@ 4.1. General setting
Grothendieck motives (for homological equivalence) form a pseudo-abelian rigid tensor category. It is built from projective smooth varieties over a field $k$, morphisms being algebraic correspondences of degree 0 modulo homological equivalence (to get a rigid tensor category, one has to invert the Lefschetz motive $h^{2}(\mathbb{P}^{1})$). In the sequel, we consider only correspondences with coefficients in $K$, the field of coefficients of the given cohomology theory.
@@
Any tensor subcategory $\mathcal{C}$} of motives cut out on products of finitely many projective smooth $k$-varieties $Y_{m}$ satisfies by definition condition P2 of §2.2. In order to also have P1 (embedding into ${\rm{Vec}}_{K}$), one needs to twist the commutativity constraint according to the Koszul rule, assuming the sign conjecture for the $Y_{m}$, i.e. that the Künneth projectors on the even part of cohomology are algebraic, so that $\mathcal{C}$} becomes even131313under Voevodsky’s smash-nilpotence conjecture, one could perform this at the level of non-integral tensor category of Chow motives; then O’Sullivan’s construction yields a $G$-scheme $\rm X$ that is a ‘fat point’ (the reduced scheme is a point), cf. [2, §2].: after twisting, the given cohomology provides a fiber functor $\omega:\,{\mathcal{C}}\to{\rm{Vec}}_{K}$.
@@
Here $\bar{\mathcal{C}}$ is the category for numerical equivalence cut out on products of $Y_{m}$’s, and $G$ (the numerical motivic Galois group) is the (reductive) tannakian group of $\bar{\mathcal{C}}$ with respect to the fiber functor given by $\bar{\omega}=\omega\sigma$, where $\sigma$ is a fixed tensor section of ${\mathcal{C}}\to\bar{\mathcal{C}}$ and $\omega$ is the fiber functor on $\mathcal{C}$} given by the selected cohomology.
According to Theorem 2.3.1, the deformation space $\rm X$} that interpolates between $\bar{\mathcal{C}}$ and $\mathcal{C}$} is the limit of a tower of fix-pointed affine quasi-homogeneous varieties with the same $G$ and same open orbit ${\rm X}_{0}=G/H$.
@@
What is $H$ in this context?
If the Künneth projectors are algebraic and char $k=0$, $H$ is the “motivated" motivic Galois group [1], the tannakian group of the tensor category ${\mathcal{C}}_{mot}$ obtained from $\mathcal{C}$} by adding formally the inverses of Lefschetz operators $L_{Y_{m}}$ (the isomorphisms $H^{i}(Y_{m})\to H^{2d-i}(Y_{m})(d-i)$ given by the (di)(d-i)th iterated cup-product with the class of an ample line bundle on the $d$-dimensional variety $Y_{m}$).
Indeed, ${\mathcal{C}}_{mot}$ is abelian, hence a target from ${\mathcal{C}}_{tann}$; and $L_{Y_{m}}$, which induces a morphism in ${\mathcal{C}}_{tann}$ with trivial kernel and cokernel, is an isomorphism - from which it is easy to conclude that ${\mathcal{C}}_{mot}={\mathcal{C}}_{tann}$. Moreover, ${\mathcal{C}}_{mot}$ is semisimple [1], hence $H$ is a reductive subgroup of $G$, and $\rm X_{0}$ is affine.
In this context, the caveat 2.1.3 takes the form the notion of a motivated class on some product $Y$ of $Y_{m}$’s may depend a priori on the whole set of $Y_{m}$’s and not only on $Y$.
@@
for $\mathcal{C}$}, is equivalent to ${\mathcal{C}}=\bar{\mathcal{C}}$, and to
@@
The deformation space $\rm X$} thus appears as an obstruction to conjecture D. Closed $G$-subvarieties in the quasi-homogeneous scheme $\rm X$ correspond to adequate equivalence relations intermediate between homological equivalence and numerical equivalence.
Even if one does not know Conjecture D for $\mathcal{C}$}, one can draw up a catalog of possibilities for $\rm X$, which often turns out to be very short. If all the items in this catalog turn out to be spherical, one deduces that there are finitely many adequate equivalence relations intermediate between homological equivalence and numerical equivalence, which can be described in terms of the geometry of $\rm X$} (as well as in terms of Asok’s filtrations 1.1.4).